Does:
5x^2/10x^-5
reduce to
0.5x^5
???

Answer by stanbon(75887)

5x^2/10x^-5
-----------
= (5/10)(x^2/x^-5)
---------
= (1/2)(x^7
---
= x^7/2
============
Cheers,
Stan H.
====================